OpenStudy (anonymous):

For any event A, 0< P(A)<1 . True False

OpenStudy (accessdenied):

With no =? That is false I can have the event of rolling a seven with a six-sided cube, but there's 0 probability of that actually happening the probability of rolling a number 1-6 is 100%, or 1. they can = the boundaries

OpenStudy (accessdenied):

\[ 0 \le P(A) \le 1 \] would be the correct inequality